Ginny's piggy bank contains $8.80 in quarters, dimes and nickels. There were 2 more than 5 times as many nickels as quarters and 4 less than twice as many dimes as quarters. How many of each kind of coin was there in the bank?
Answer by TimothyLamb(4379)

---
5x + 10y + 25z = 880
x = 5z + 2
y = 2z - 4
---
put the system of linear equations into standard form
---
5x + 10y + 25z = 880
x - 5z = 2
y - 2z = -4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
x= 67
y= 22
z= 13
---
answer:
nickels = 67
dimes = 22
quarter = 13
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php